OpenStudy (amistre64):

Proof by Contradiction. The sum of an irrational number and a rational number is irrational. So how do we symbolically express an irrational number?

OpenStudy (amistre64):

id say 1/0 is irrational :)

OpenStudy (jamesj):

What's the definition of a rational number?

OpenStudy (amistre64):

p/q such that p and q are ints and q not 0

OpenStudy (jamesj):

Right

OpenStudy (jamesj):

So let x be an irrational number and y a rational number. Let us suppose that in fact x + y is rational, then ....

OpenStudy (jamesj):

x + y = ....

OpenStudy (amistre64):

good, ill try to work on that angle :)

OpenStudy (amistre64):

Let m be any irrational number, and let n be any rational number such that; m+n = p/q, q not 0. m + \(\cfrac{p_1}{q_1}\) = \(\cfrac{p}{q}\) m = \(\cfrac{p}{q}\) - \(\cfrac{p_1}{q_1}\) m = \(\cfrac{p q_1 - qp_1}{q\ q_1}\) = False maybe?

OpenStudy (jamesj):

Yes, because by hypothesis m is irrational, yet now you have shown is rational. That is a cotradiction. Hence it cannot be the case that if m is irrational and n is rational then m + n is rational Therefore: if m is irrational and n is rational then m + n is irrational.

OpenStudy (amistre64):

thanx :)

OpenStudy (jamesj):

make that contradiction.

OpenStudy (jamesj):

Proof by contradiction is a very important tool in mathematics. In general, suppose you want to show that A => B Logically this is exactly equivalent to not ( A and not B) Hence if we can show that A and not B is false, then it must indeed be the case that not (A and not B) is true. I.e., that A => B.

OpenStudy (jamesj):

That's exactly what we've done here with A = m is irrational, n is rational B = m + n is irrational You started with the statement A, and the statement not B, and showed that A and not B can't be true.